Help!!! I got the first part, but am having trouble with the rest.
Lauren attended her 10th high school reunion. Of the 74 people present, she observed that 37 were married and in their first marriage, 23 were married and in their second marriage, and 14 were currently single. (to three decimal places.)
(a) If one person is selected at random from the group, find the probability that it is a single person.
For a I got a correct answer of 0.189, but I can't figure out how to get the rest :(
(b) If two people are selected at random, find the probability that both are in their second marriage.
(c) If three people are selected at random, find the probability that two are in their first marriage and one is single.

Answer by stanbon(75887)

Lauren attended her 10th high school reunion.
---------------------------------------------------
Of the 74 people present, she observed that
37 were married and in their first marriage,
23 were married and in their second marriage, and
14 were currently single. (to three decimal places.)
-----------------------------
(a) If one person is selected at random from the group, find the probability that it is a single person.
For a I got a correct answer of 0.189, but I can't figure out how to get the rest :(
(b) If two people are selected at random, find the probability that both are in their second marriage.
---
Ans: 23C2/74C2 = 253/2701 = 0.094
----------------------------
(c) If three people are selected at random, find the probability that two are in their first marriage and one is single.
Ans: [37C2*14]/74C3 = 9324/64824 = 0.144
=============
Cheers,
Stan H.